OpenStudy (anonymous):

Subtract and reduce the following: 1/9 - 1/10

OpenStudy (anonymous):

is the answer 1

OpenStudy (anonymous):

They need common denominators. 10 / 90 - 9 / 90 = 1 / 90.

OpenStudy (anonymous):

why is it 90? u multiply?

OpenStudy (anonymous):

Divide and reduce: 7/12 / 12 /16=

OpenStudy (anonymous):

is the correct answer 84/160

OpenStudy (anonymous):

Add and reduce 4 3/7 +2 5/9 + 4 + 2/3 =

OpenStudy (anonymous):

i got 11 1/3 is that correct?